OpenStudy (anonymous):

Arithmetic Sequence A man puts $300 into a savings account at the end of each year for 15 yrs. If his money draws 3% simple interest, how much will he have in the account at the end of 15 yrs?

OpenStudy (anonymous):

M = C + I I = Cdt M = C + Cdt M = 300 + (300)(0.03)(15) M = 300 + 135 M= 435$

OpenStudy (anonymous):

D being the interest in decimal form, C being the initial deposit or capital, t being the time in years and M being the total money in the account.

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

@GodlyMaths the result is not 426? just making sure

OpenStudy (anonymous):

No, I'm quite confuse